OpenStudy (anonymous):

can any1 help me with physics?

OpenStudy (anonymous):

electro magnet waves, and how to solve problems.

OpenStudy (unklerhaukus):

ok

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

well the thing is i have approx 2 months left for a physics exam,

OpenStudy (anonymous):

and i need to know the hole book, it is like 300 pages, of problems, examples etc etc

OpenStudy (anonymous):

i do understand some of the chapters, but some of them like the problems

OpenStudy (unklerhaukus):

i see

OpenStudy (anonymous):

the intensite level in a street is 70 db, what is the sound time rate of the energu entering a room trough a windows of area 3m^2 facing the street

OpenStudy (anonymous):

this it the "sound" chapter

OpenStudy (unklerhaukus):

the window acts as a surface source, find the intensity of sound at the window,

OpenStudy (unklerhaukus):

then use the area of the window to find the total intensity in the room due to the window

OpenStudy (anonymous):

70 = 10log I/10^-12 = I / 10^-12 = 7

OpenStudy (anonymous):

something like that lol, buts is was to hard lol

OpenStudy (anonymous):

sound rate of energy = 10^-5 watt/m2 x 3m^2 = 3x10^-5 watt....

OpenStudy (anonymous):

no one that is good at this things? ill even pay trough paypal if ur good lol

OpenStudy (anonymous):

then we have ac circuits, diffraction gratings, modern physics etc etc

OpenStudy (vincent-lyon.fr):

"the intensite level in a street is 70 db, what is the sound time rate of the energu entering a room trough a windows of area 3m^2 facing the street" 70dB is the acoustic level. It is an arbitrary logarithmic scale for sound surface-power. So sound intensity is \(I=10^{-7}\: Wm^{-2}\) And power coming in through the window is \(P=3.10^{-5}W\) Your answer was ok :)